OpenStudy (anonymous):

Help can someone tell me in as much detail as possible how to do this?!!! We are looking for the rate of change and make sure to simplify the answer.

OpenStudy (johnweldon1993):

Question?

OpenStudy (anonymous):

Here is the problem

OpenStudy (johnweldon1993):

So first, notice that 'rate of change' is simply a way of saying 'slope' So we want the slope of the line that passes through those 2 points right?

OpenStudy (anonymous):

Yes

OpenStudy (johnweldon1993):

Great, so all we need to do is use the slope formula \[\large m = \frac{y_2 - y_1}{x_2 - x_1}\] We have 2 points...we can write them as (0, 18) as (x1 and y1) and (16, 8) as (x2 and y2) So all we need to do is plug them into that formula and simplify

OpenStudy (anonymous):

@johnweldon1993 ok

OpenStudy (anonymous):

Can you show me how that would look? Sorry:/

OpenStudy (johnweldon1993):

No that's fine, so the formula again \[\large m = \frac{y_2 - y_1}{x_2 - x_1}\] We know that x1 = 0 x2 = 16 y1 = 18 y2 = 8 SO we have \[\large m = \frac{8 - 18}{16 - 0}\] \[\large m = \frac{-10}{16}\] \[\large m = \frac{-5}{8}\]

OpenStudy (anonymous):

Thank you thank you thank you!!!